Jennifer's age is 3 less than 8 times of David's age. 21 years later, the ratio of Jennifer's age to David's age will be 1:2. Find Jennifer's current age.
Answer by lwsshak3(11628)

Jennifer's age is 3 less than 8 times of David's age. 21 years later, the ratio of Jennifer's age to David's age will be 1:2. Find Jennifer's current age.
(I will assume you meant the ratio to be 2:1 instead of 1:2)
***
let x=david's age
8x-3=jennifer's age
..
21 yrs later:
x+21=david's age
8x-3+21=8x+18=jennifer's age
(8x+18)/(x+21)=2/1
(8x+18)=2(x+21)
8x+18=2x+42
6x=24
x=4
8x-3=29
jennifer's age=29